# Performance Optimization of High-Conflict Transactions within the Hyperledger Fabric Blockchain

### v1

@@ IV-A OEMVCC
OEMVCC takes advantage of the latest versions of HLF, which can incorporate VSCC on the GW, invalidating transactions before dispatching them to the ordering service and notifying the client earlier in the transaction flow, as presented in Figure 3 (in OEMVCC). Performing VSCC directly after the execution phase on the GW guarantees that valid MVCC transactions on orderers will successfully pass VSCC on the validation phase and eventually change the version of their keys upon block commitment.
@@
The caching mechanism within the ordering service should account for the logic of the underlying consensus algorithm. Currently, the default consensus algorithm is Raft, consisting of a leader and its followers. It regularly elects a leader orderer who orchestrates the ordering of transactions into blocks for its leadership duration. Followers forward incoming transactions to their leader, ensuring strong consistency between their log entries and those of the leader. In OEMVCC, the leader also performs MVCC, implying that the caching mechanism should incorporate strong consistency in its entries to accommodate elections, as a follower can transition into a leader at any time and undertake the ordering and MVCC. Thus, we implement a distributed cache coordinated by the consensus protocol to maintain this strong consistency and incorporate MVCC in the ordering service.
The behavior of peers during the validation phase also depends on whether OEMVCC or OEMVCC-EA is employed, as outlined in Algorithm 3. Contrary to the EOV-OG, where transactions sequentially undergo VSCC and MVCC, upon receiving a block, peers in OEMVCC, depicted in Figure 4 (in OEMVCC), bypass the VSCC and MVCC validation procedures, as the ordering service has already designated all transactions within the block valid or invalid. Thus, the peer can directly commit the block to the ledger (line 14). Finally, we extend the GW to support receiving and handling notifications by the ordering service to notify the client regarding invalid transactions.
@@
4:         if $ea$ then
5:              for $tx$ in $block$ do
6:                  if $tx.mvcc=0$ then
7:                       continue
8:                  end if
9:                  for $key$ in $tx.wset$ do
10:                       $peer\rightarrow cache[key]\leftarrow 0$
11:                  end for
12:              end for
13:         end if
14:         Commit the block to the ledger
15:     end while
16:end function
@@ IV-B OEMVCC-EA
OEMVCC-EA extends upon OEMVCC, aiming to increase transaction throughput while conserving network and computational resources by dropping invalid transactions earlier in the transaction flow. Figure 2 (in OEMVCC-EA) and Algorithm 1 ($ea$ is true) describe how the ordering service incorporates OEMVCC-EA. Similarly to OEMVCC, in OEMVCC-EA, the GW performs VSCC, and the ordering service is MVCC. OEMVCC-EA reduces block sizes in the ordering phase by discarding invalid transactions after MVCC (lines 8-10), ensuring blocks contain solely valid transactions. Consequently, the ordering service proactively notifies the client about invalid transactions via the GW (line 11) and informs the peers about the keys in the write sets of valid transactions (lines 12-18) to optimize their resource utilization in the execution phase.
@@
7:                  if $peer\_cache[tx.key]=1$ then
@@
Endorsing peers utilize the write sets of valid transactions to avoid simulating transactions containing these keys since the resulting read-write sets will eventually become outdated, and the transaction will ultimately fail, as described in Figure 3 (in OEMVCC-EA) and Algorithm 4 ($ea$ is true) (lines 3 - 14). Finally, as the peers receive blocks containing exclusively valid transactions, they can directly commit the included transactions, as illustrated in Figure 4 (in OEMVCC-EA) and described in Algorithm 3. Once a transaction is committed to the ledger by a peer, the peer can remove the keys included on the write of the validated transaction (lines 9-11), enabling it to endorse upcoming transactions that involve these keys.

### v2

@@ IV-A OEMVCC
OEMVCC takes advantage of the latest versions of HLF, which can incorporate VSCC on the GW, invalidating transactions before dispatching them to the ordering service and notifying the client earlier in the transaction flow, as presented in Figure 3 (in OEMVCC). Performing VSCC directly after the execution phase on the GW increases the probability that valid MVCC transactions on orderers will successfully pass VSCC on the validation phase and eventually change the version of their keys upon block commitment. Peers must still perform VSCC on the validation phase to avoid malicious clients that might forge the results of the execution phase and submit the transaction directly to the ordering service.
@@
The caching mechanism within the ordering service should account for the logic of the underlying consensus algorithm. Raft regularly elects a leader orderer who orchestrates the ordering of transactions into blocks for its leadership duration. Followers forward incoming transactions to their leader, ensuring strong consistency between their log entries and those of the leader. In OEMVCC, the leader also performs MVCC, indicating that the caching mechanism should incorporate strong consistency in its entries to accommodate elections, as a follower can transition into a leader at any time and undertake the ordering and MVCC. Each orderer maintains one write-through cache for each channel it belongs to, ensuring privacy and isolation between channels. The leader updates the caches of its followers after it orders a transaction into a block and before ordering subsequent transactions. When the leader crashes or when a new leader is elected before the caches are updated, the new leader is responsible for updating the caches according to its log entries or the incomplete block received. Finally, if the caches become full, the orderers transfer the oldest entries to persistent storage.
The behavior of peers during the validation phase also depends on whether OEMVCC or OEMVCC-EA is employed, as outlined in Algorithm 3. Contrary to the EOV-OG, where transactions sequentially undergo VSCC and MVCC, upon receiving a block, peers in OEMVCC, depicted in Figure 4 (in OEMVCC), solely perform VSCC to detect transactions submitted by malicious clients, which bypass VSCC on the GW (lines 5-9). After VSCC, the peer can commit the block directly to the ledger (line 16), avoiding MVCC. Finally, we extend the GW to support receiving and handling notifications by the ordering service to notify the client regarding invalid transactions.
@@
4:         for $tx$ in $block$ do
5:              if $not\ ea$ then
6:                  $tx.vscc\leftarrow VSCC(tx)$
7:                  if $tx.vscc=0\text{ OR }tx.mvcc=0$ then
8:                       continue
9:                  end if
10:              else
11:                  for $key$ in $tx.wset$ do
12:                       $peer\rightarrow cache[key]\leftarrow 0$
13:                  end for
14:              end if
15:         end for
16:         Commit the block to the ledger
17:     end while
18:end function
@@ IV-B OEMVCC-EA
OEMVCC-EA extends upon OEMVCC, aiming to increase transaction throughput while conserving network and computational resources by dropping invalid transactions earlier in the transaction flow. Figure 2 (in OEMVCC-EA) and Algorithm 1 ($ea$ is true) describe how the ordering service incorporates OEMVCC-EA. OEMVCC-EA assumes no malicious clients exist on the network and performs VSCC exclusively on the GW. Since the GW and the ordering service detect invalid VSCC and MVCC transactions, block sizes can be reduced by discarding invalid transactions after the ordering phase (lines 8-10), creating blocks that contain solely valid transactions, and notifying the clients earlier through the GW (line 11). The ordering service proactively informs the peers about the keys in the write sets of valid transactions to optimize their execution and validation phases (lines 12-18).
@@
7:                  if $peer\rightarrow cache[tx.key]=1$ then
@@
Endorsing peers utilize the write sets of valid transactions to avoid simulating transactions containing these keys since the resulting read-write sets will eventually become outdated, and the transaction will ultimately fail, as described in Figure 3 (in OEMVCC-EA) and Algorithm 4 ($ea$ is true) (lines 3 - 14). Finally, as the peers receive blocks containing exclusively valid transactions, they can directly commit the included transactions, as illustrated in Figure 4 (in OEMVCC-EA) and described in Algorithm 3. Once a transaction is committed to the ledger by a peer, the peer can remove the keys included on the write of the validated transaction (lines 11-13), enabling it to endorse upcoming transactions that involve these keys.